OpenStudy (anonymous):

f(x)=x^2-f and g(x)= 2x+1 Find f(g(x)) help!

OpenStudy (anonymous):

\[f(x)=x^2-4 \]

OpenStudy (anonymous):

\[g(x)=2x+1\]

OpenStudy (anonymous):

Well ill try this: \[f(g(x))=f(2x+1)=(2x+1)^2-4=4x^2+4x+1-4=4x^2+4x-3\]

OpenStudy (anonymous):

Woah... What?

OpenStudy (anonymous):

I think I understand where you got the f(2x+1)? But I'm not sure...

OpenStudy (anonymous):

Basically when you say \(f(g(x))\), your referring to the operation of wherever x was in f(x), your putting \(g(x)\) instead!

OpenStudy (anonymous):

So you would have \[f(g(x))=f(2x+1)=(2x+1)^2-4\] Because before, \(f(x)=x^2-4\) But now im putting g(x) where x was.

OpenStudy (anonymous):

Okay, I see that. Now what for the next problem it wants me to find 2f(3)+3g(-2) I don't get it.

OpenStudy (anonymous):

OOOOH Ouch that's okay. So: \[\eqalign{ &f(x)=x^2-4 \\ &g(x)=2x+1 \\ }\] Still right?

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

Would I just put 3 in for f(x) then it would make: \[2=3^2-4\]